OpenStudy (anonymous):

What is the equation of a line passing through (3, 4) and having a slope of -5?

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

m= slope, y1= y coordinate, x1= x coordinate

OpenStudy (anonymous):

take the information you are given and plug in the numbers into that and solve for y and that will give you your equation :)

OpenStudy (anonymous):

I need the answer im on a test -.-

OpenStudy (anonymous):

i no yur not suppose to but please

OpenStudy (anonymous):

y - 4 = -5(x-3) y-4 =-5x + 15 y = -5x +19

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):

sure thing good luck on the rest of your test!